# The Reward Model Selection Crisis in Personalized Alignment

### v1

@@ Abstract
Personalized alignment from preference data has focused primarily on improving reward model (RM) accuracy, with the implicit assumption that better preference ranking translates to better personalized behavior. However, in deployment, computational constraints necessitate inference-time adaptation via reward-guided decoding (RGD) rather than per-user policy fine-tuning. This creates a critical but overlooked requirement: reward models must not only rank preferences accurately but also effectively guide token-level generation decisions. We demonstrate that standard RM accuracy fails catastrophically as a selection criterion for deployment-ready personalized alignment. Through systematic evaluation across three datasets, we introduce policy accuracy—a metric quantifying whether RGD scoring functions correctly discriminate between preferred and dispreferred responses—and show that RM accuracy correlates only weakly with this discrimination ability (Kendall’s $\tau$ = 0.08–0.31). More critically, we introduce Pref-LaMP, the first personalized alignment benchmark with ground-truth user completions, enabling direct behavioral evaluation without circular reward-based metrics. On Pref-LaMP, we expose a complete decoupling between discrimination and generation: methods with 20-point RM accuracy differences produce almost identical output quality, and even methods achieving high discrimination fail to generate behaviorally aligned responses. Furthermore, simple in-context learning (ICL) dominates all reward-guided methods for models $\geq$3B parameters, achieving 3–5 point ROUGE-1 gains over the best reward method at 7B scale. These findings reveal that the field has been optimizing for proxy metrics that don’t predict deployment performance, and that current personalized alignment methods fundamentally fail to operationalize preferences into behavioral adaptation under realistic deployment constraints.
@@
Fady Rezk1,2, Yuangang Pan2, Chuan-Sheng Foo2, Xun Xu2, Nancy Chen2, Henry Gouk1, Timothy Hospedales1,3 1 School of Informatics, University of Edinburgh, UK 2 Agency for Science, Technology and Research (A*STAR), Singapore 3 Samsung AI Research Center, Cambridge, UK
@@ 1 Introduction
Current personalized alignment research has converged on a common paradigm: collect user-specific preference data (pairwise comparisons, ratings, or implicit signals), train personalized reward models to capture individual preferences, and assume that better reward models naturally translates to better policies (Bose et al., 2025; Chen et al., 2025a; Shenfeld et al., 2025; Li et al., 2024; Poddar et al., 2024). The last assumption is known to break as reasoned by Goodhart’s law El-Mhamdi and Hoang (2024). Unlike standard RLHF, there are no downstream benchmark that measures policy performance such as MMLU Hendrycks et al. (2021) and GSM8k Cobbe et al. (2021).
1.1 The Deployment Reality and End-to-End Perspective
Per-user policy fine-tuning using reward model artefacts via RL is computationally infeasible at scale. For 1000 user groups, RL-based personalization requires dynamic adapter management, RL instability mitigation, and orders of magnitude more compute than inference-time alternatives. The only scalable path is inference-time adaptation through reward-guided decoding (RGD) Khanov et al. (2024), maintaining a single base policy while using personalized rewards to guide generation. Another option is or Best-of-N sampling Ichihara et al. (2025) but the high latency of BoN makes it unfit for personalization text generation.
@@ 1.1 The Deployment Reality and End-to-End Perspective
A personalized alignment method must specify not only how preferences are modeled, but how they are operationalized into behavioral adaptation. Papers proposing reward models are responsible for demonstrating that improved RM accuracy translates to improved generation behavior.
Current evaluation ignores this responsibility, treating reward modeling and policy adaptation as isolated steps. This obscures whether methods actually achieve their objective: making models generate responses aligned with user preferences.
1.2 Our Investigation
We adopt an end-to-end perspective, studying the complete chain from preference modeling to generation behavior. We ask three questions: (1) Does RM accuracy predict discrimination under RGD? (2) Does discrimination predict generation quality? (3) How do reward methods compare to simpler baselines?
To answer these, we introduce policy accuracy—measuring whether RGD scoring assigns higher scores to preferred responses—and Pref-LaMP, a benchmark with ground-truth user completions enabling direct behavioral evaluation.
Our findings reveal a fundamental selection crisis: methods with 20-point RM accuracy differences produce identical generation quality, and correlations between RM accuracy and downstream performance are negligibly weak (Kendall’s $\tau$ = 0.08-0.31). Practitioners cannot reliably choose deployment-ready methods because standard metrics don’t predict actual performance.
1.3 Key Findings
Finding 1: RM accuracy doesn’t predict discrimination (Kendall’s $\tau$ = 0.08–0.31). Methods with 20-point RM accuracy differences achieve nearly identical policy performance.
Finding 2: Discrimination doesn’t predict generation. On Pref-LaMP, methods cluster tightly in generation quality despite dramatic variance in RM and policy accuracy.
Finding 3: ICL dominates at scale. At 3B-7B parameters, ICL-RAG outperforms all reward methods by 3-5 ROUGE-1 points.
Implications: For practitioners, use ICL-RAG at scale. For researchers, RM accuracy is broken—evaluate on ground-truth completions. The field needs behavioral benchmarks like Pref-LaMP, analogous to GSM8K/MMLU for general RLHF.
1.4 Contributions
We (1) formalize the RM selection problem and introduce policy accuracy for discrimination measurement, (2) contribute Pref-LaMP—the first benchmark with ground-truth user completions, (3) demonstrate RM accuracy fails as a selection criterion across three datasets and four scales, and (4) provide actionable recommendations for practitioners and researchers.
@@ 2.1 Personalized Alignment
Recent work has focused on learning user-specific reward models or policies for alignment under limited supervision (e.g., PAL, PReF, LoRE, P-DPO, VPL) (Chen et al., 2025a; Poddar et al., 2024; Bose et al., 2025; Shenfeld et al., 2025; Li et al., 2024). These approaches largely target reward-modeling accuracy or relative preference ranking as proxies for personalization quality Chen et al. (2025a); Bose et al. (2025); Shenfeld et al. (2025). However, such metrics often fail to capture whether a policy’s generated responses truly reflect an individual user’s preferences, especially under resource-constrained settings or whether it is an unintended overoptimization (Pan et al., 2022).
2.2 Multi-Objective Alignment
@@
Multi-objective alignment (MOA) addresses the challenge of optimizing language models across multiple known and predefined reward dimensions simultaneously. Unlike personalized alignment, where the goal is to learn individual user preferences under limited supervision, MOA assumes access to distinct reward models for each objective dimension (e.g., helpfulness, harmlessness, factuality) and focuses on finding optimal policy trade-offs among these objectives. Prior work has explored weighted reward optimization (Zhou et al., 2024), model merging (Jang et al., 2024; Rame et al., 2023), auxiliary correction models (Ji et al., 2024; Yang et al., 2024a), test-time reward-guided decoding Chen et al. (2025b) among other methods (Yang et al., 2024b). Despite the terminological overlap with ”personalization” in some methods, the key distinction is that MOA operates over known objective dimensions with pre-trained reward models, whereas personalized alignment learns latent user-specific preferences from sparse behavioral data across dataset specific dimensions (e.g., style, length, political values, rhetoric, culture sensitivities, etc).
2.3 Evaluation Challenges in RLHF
@@
Existing evaluation practices in RLHF and personalized alignment rely on proxy metrics such as reward-model scores which are susceptible to reward hacking and circularity (Tien et al., 2023; Pan et al., 2022). These methods assess optimization success rather than behavioral fidelity Wen et al. (2025); Gao et al. (2023); El-Mhamdi and Hoang (2024). In contrast, our work introduces a framework for direct behavioral evaluation, measuring whether generated responses match user-provided completions (Section 5). Extended discussion of related evaluation pathologies appears in Appendix A.
Our work addresses these limitations by introducing direct behavioral evaluation on ground-truth user completions, measuring actual generation quality rather than relying on proxy metrics. Extended discussion of evaluation pathologies appears in Appendix A.
2.4 Inference-Time Alignment
@@
Reward-guided decoding (ARGS) Khanov et al. (2024) and Best-of-N sampling Ichihara et al. (2025) enable policy steering without fine-tuning, making them computationally attractive for personalization. Recent work has explored their effectiveness Wu (2025), but evaluation remains limited to reward-based metrics. Our work is the first to systematically evaluate these methods for personalized alignment with ground-truth behavioral assessment, revealing fundamental limitations in their ability to operationalize user preferences.
@@ 3.1 Personalized Preference Learning
Consider a preference dataset $\mathcal{D}=\{(u_{i},x_{i},y_{i}^{(c)},y_{i}^{(r)})\}_{i=1}^{N}$, where $u_{i}\in U=\{1..K\}$ is a user identifier, $x_{i}$ is a prompt, and $y_{i}^{(c)},y_{i}^{(r)}$ are chosen and rejected completions. We split users into training $\mathcal{U}_{\mathrm{train}}$ (learning shared structure) and adaptation $\mathcal{U}_{\mathrm{adapt}}$ (few-shot personalization at deployment). For user $k$, define $\mathcal{D}_{k}=\{(x_{i},y_{i}^{(c)},y_{i}^{(r)}):u_{i}=k\}$.
Personalized Reward Modeling. The reward model conditions on user identity: $r_{\theta,z_{k}}(y\mid x)$, decomposing into shared parameters $\theta$ (general preference structure) and user-specific parameters $z_{k}$ (individual preferences). Training on $\mathcal{U}_{\mathrm{train}}$ learns both $\theta$ and $\{z_{k}\}_{k\in\mathcal{U}_{\mathrm{train}}}$. At deployment, for new user $k\in\mathcal{U}_{\mathrm{adapt}}$ with few-shot data $\mathcal{D}_{k}^{\mathrm{few-shot}}$, we adapt:
$z_{k}=\mathcal{A}_{R}\big(\mathcal{D}_{k}^{\mathrm{few-shot}};\theta\big)$
where $\mathcal{A}_{R}$ is the adaptation algorithm.
@@ 3.2 Deployment via Reward-Guided Decoding
Given computational constraints prohibiting per-user policy fine-tuning, we deploy personalized alignment through inference-time guidance. We consider two methods:
Reward-Guided Decoding (ARGS). At each generation step $t$, ARGS scores the top-$k$ tokens $V^{(k)}_{t}$ as:
@@
$\quad+\lambda\cdot r_{\theta,z_{k}}(v\mid x,y_{<t})$
@@ 3.3 The Evaluation Gap
Standard practice evaluates personalization methods by reward model accuracy:
@@
For user $k$ with evaluation set $\mathcal{D}_{k}^{\text{eval}}=\{(x_{i},y_{i}^{(c)},y_{i}^{(r)})\}$:
$\frac{1}{|\mathcal{D}_{k}^{\text{eval}}|}\sum_{\mathcal{D}_{k}^{\text{eval}}}\mathbb{I}\Big[r_{\theta,z_{k}}(y_{i}^{(c)}\mid x_{i})>r_{\theta,z_{k}}(y_{i}^{(r)}\mid x_{i})\Big]$ (2)
This measures pairwise ranking on complete responses. However, deployment via RGD requires token-level scoring. A reward model achieving high sequence-level ranking accuracy may provide poor token-level guidance, yet standard evaluation cannot detect this mismatch.
4 Policy Accuracy: Measuring Discrimination Under RGD
To bridge the evaluation gap, we introduce a metric quantifying whether the RGD scoring function—not just the reward model in isolation—can discriminate preferred from dispreferred responses.
@@ 4 Policy Accuracy: Measuring Discrimination Under RGD
Let $s_{\pi}:\mathcal{Y}\times\mathcal{X}\to\mathbb{R}$ be the policy scoring function used at generation time. The policy accuracy for user $k$ is:
$\frac{1}{|\mathcal{D}_{k}^{\text{eval}}|}\sum_{\mathcal{D}_{k}^{\text{eval}}}\mathbb{I}\Big[s_{\pi}(y_{i}^{(c)}\mid x_{i})>s_{\pi}(y_{i}^{(r)}\mid x_{i})\Big]$ (3)
We instantiate $s_{\pi}$ with three scoring functions, each revealing different aspects of the personalization pipeline:
Prior Policy. The base policy’s length-normalized log-likelihood, quantifying zero-shot discrimination ability:
$\frac{1}{|y|}\sum_{t=1}^{|y|}\log\pi(y_{t}\mid x,y_{<t})$ (4)
Global RM Posterior. RGD scoring with a non-personalized reward model:
$\sum_{t=1}^{|y|}\Big[\log\pi(y_{t}\mid x,y_{<t})+\lambda\cdot r_{\theta}(y_{t}\mid x,y_{<t})\Big]$ (5)
Personalized Posterior. RGD scoring with personalized reward model:
$\sum_{t=1}^{|y|}\Big[\log\pi(y_{t}\mid x,y_{<t})$ $\quad+\lambda\cdot r_{\theta,z_{k}}(y_{t}\mid x,y_{<t})\Big]$ (6)
Comparing these accuracies reveals: (1) whether reward guidance improves discrimination over the base policy, and (2) whether personalization provides gains beyond global reward models.
Critical Limitation. Policy accuracy measures discrimination ability—whether the scoring function prefers chosen responses—not generation quality. High policy accuracy provides no guarantee the policy will generate outputs that behaviorally align with user preferences. A method might correctly rank static responses while producing generations that differ substantially from what users would actually write. This motivates our ground-truth behavioral evaluation in Section 5.

### v2

@@ Abstract
Personalized alignment from preference data has focused primarily on improving personal reward model (RM) accuracy, with the implicit assumption that better preference ranking translates to better personalized behavior. However, in deployment, computational constraints necessitate inference-time adaptation such as reward-guided decoding (RGD) rather than per-user policy fine-tuning. This creates a critical but overlooked requirement: reward models must not only rank preferences accurately but also effectively guide generation. We demonstrate that standard RM accuracy fails catastrophically as a selection criterion for deployment-ready personalized rewards. We introduce policy accuracy—a metric quantifying whether RGD-adapted LLMs correctly discriminate between preferred and dispreferred responses—and show that upstream RM accuracy correlates only weakly with downstream policy accuracy (Kendall’s $\tau$ = 0.08–0.31). More critically, we introduce Pref-LaMP, the first personalized alignment benchmark with ground-truth user completions, enabling direct behavioural evaluation. On Pref-LaMP, we expose a complete decoupling between discriminative ranking and generation metrics: methods with 20-point RM accuracy differences produce almost identical output quality, and methods with high ranking accuracy can fail to generate behaviorally aligned responses. These findings reveal that the field has been optimizing for proxy metrics that do not predict deployment performance, and that current personalized alignment methods fail to operationalize preferences into behavioral adaptation under realistic deployment constraints. In contrast, we find simple in-context learning (ICL) to be highly effective - dominating all reward-guided methods for models $\geq$3B parameters, achieving $\sim$3 point ROUGE-1 gains over the best reward method at 7B scale.
@@
Fady Rezk1,2, Yuangang Pan2, Chuan-Sheng Foo2, Xun Xu2, Nancy F. Chen2, Henry Gouk1, Timothy Hospedales1,3 1 School of Informatics, University of Edinburgh, UK 2 Agency for Science, Technology and Research (A*STAR), Singapore 3 Samsung AI Research Center, Cambridge, UK
@@ 1 Introduction
Current personalized alignment research has converged on a common paradigm: collect user-specific preference data (pairwise comparisons), train personalized ranking/reward models to capture individual preferences, and assume that better reward models naturally translates to better policies (Bose et al., 2025; Chen et al., 2025a; Shenfeld et al., 2025; Li et al., 2024; Poddar et al., 2024). The last assumption is likely to break as suggested by Goodhart’s law El-Mhamdi and Hoang (2024). Unlike standard RLHF, personalized alignment lacks downstream benchmarks that measures policy performance such as MMLU Hendrycks et al. (2021) and GSM8k Cobbe et al. (2021).
Practical Deployment and the End-to-End Perspective  Per-user policy fine-tuning using personal rewards via RL is computationally infeasible at scale. RL-based personalization requires per-user dynamic adapter management, RL instability mitigation, and orders of magnitude more compute than inference-time alternatives. One key scalable deployment path is inference-time adaptation through reward-guided decoding (RGD) Khanov et al. (2024), maintaining a single base policy while using personalized rewards to guide generation. Another option is or Best-of-N sampling Ichihara et al. (2025) but the high latency of BoN makes it unfit for personalization text generation.
@@
A personalized alignment method must specify not only how preferences are modeled, but how they are operationalized into behavioral adaptation.
A direct corollary of this is that papers proposing reward models are responsible for evaluating if improved RM accuracy translates to improved generation. Current evaluations ignore this responsibility, treating reward modeling and policy adaptation as independent. This obscures whether methods actually achieve their objective: making models generate responses aligned with user preferences.
Our Investigation  We adopt an end-to-end perspective, studying the complete chain from preference modeling to generation behavior. We ask three questions: (1) Does RM accuracy predict policy accuracy under RGD? (2) Does policy accuracy predict generation quality? (3) How do reward-based alignment methods compare to simpler baselines? To answer these, we introduce (A) policy accuracy, measuring whether RGD scoring assigns higher scores to preferred responses, and (B) Pref-LaMP, a preference learning benchmark with ground-truth user completions enabling direct behavioral evaluation.
Key Findings  Our findings reveal a fundamental selection crisis: practitioners cannot reliably choose deployment-ready methods because standard metrics do not predict actual performance.
Finding 1: Upstream RM accuracy does not predict downstream policy accuracy under RGD (Kendall’s $\tau$ = 0.08–0.31). Methods with 20-point RM accuracy differences achieve nearly identical policy performance.
Finding 2: Response ranking quality does not predict response generation quality. On Pref-LaMP, methods with similar generation quality vary dramatically in RM and policy accuracy.
Finding 3: ICL dominates at scale. At 7B parameters, ICL-RAG, with RAG selected preference demonstrations Salemi et al. (2024), outperforms best personalized reward model by $\sim$3 ROUGE-1 points.
Implications: For practitioners, use simple ICL-RAG in preference to published personal reward methods. For researchers, take an end-to-end perspective: co-design and co-evaluate reward personalization with policy adaptation strategies and evaluate generation quality as well as ranking. Use Pref-LaMP and develop more benchmarks with ground-truth completions, analogous to GSM8K/MMLU for general RLHF.
Contributions  To summarize, we (1) contribute Pref-LaMP—the first benchmark with ground-truth user completions, (2) demonstrate the standard RM accuracy metric fails as a selection criterion across three datasets and four scales, (3) demonstrate that a simple ICL baseline outperforms published personal alignment work in end-to-end adaptation, and (4) provide actionable recommendations for practitioners and researchers.
@@ 2 Related Work
Personalized Alignment  Recent work has focused on learning user-specific reward models or policies for alignment under limited supervision (e.g., PAL, PReF, LoRE, P-DPO, VPL) (Chen et al., 2025a; Poddar et al., 2024; Bose et al., 2025; Shenfeld et al., 2025; Li et al., 2024). These approaches largely target reward-modeling accuracy (e.g relative preference ranking) as proxies for personalization quality Chen et al. (2025a); Bose et al. (2025); Shenfeld et al. (2025). However, such metrics often fail to capture (i) whether RM adaptation translates to downstream policy adaptation, and (ii) whether, under realistic resource-constrained settings, a personalized policy is able to go beyond response ranking and actually generate responses reflective of a user’s preferences. This evaluation limitation leaves open the null hypothesis that prior personal alignment results, measured by RM accuracy, are due to unintended overoptimization, also known as reward-hacking (Pan et al., 2022).
@@
Multi-Objective Alignment  Multi-objective alignment (MOA) addresses the challenge of optimizing language models across multiple known and predefined reward dimensions simultaneously. Unlike personalized alignment, where the goal is to learn individual user preferences under limited supervision, MOA assumes access to distinct reward models for each objective dimension (e.g., helpfulness, harmlessness, factuality) and focuses on finding optimal policy trade-offs among these objectives. Prior work has explored weighted reward optimization (Zhou et al., 2024), model merging (Jang et al., 2024; Rame et al., 2023), auxiliary correction models (Ji et al., 2024; Yang et al., 2024a), test-time reward-guided decoding Chen et al. (2025b) among other methods (Yang et al., 2024b).
@@
Evaluation Challenges in RLHF  Existing evaluation practices in RLHF and personalized alignment rely on proxy metrics such as reward-model scores which are susceptible to reward hacking and circularity (Tien et al., 2023; Pan et al., 2022). These methods assess optimization success rather than behavioral quality Wen et al. (2025); Gao et al. (2023); El-Mhamdi and Hoang (2024). In contrast, our work introduces a framework for direct behavioral evaluation, measuring whether generated responses match user-provided completions (Section 5). Extended discussion of related evaluation pathologies appears in Appendix A. Our work addresses these limitations by introducing direct behavioral evaluation on ground-truth user completions, measuring actual generation quality rather than relying on proxy metrics.
@@
Inference-Time Alignment  Reward-guided decoding Khanov et al. (2024) and Best-of-N sampling Ichihara et al. (2025) enable policy steering without fine-tuning, making them computationally attractive for personalization. Recent work has explored their effectiveness Wu (2025), but standard personal alignment evaluation remains limited to reward-based metrics. Our work is the first to systematically evaluate test-time alignment (reward-guided decoding in particular) for personalized alignment with ground-truth behavioral assessment, revealing fundamental limitations in their ability to operationalize user preferences.
@@ 3.1 Personalized Preference Learning
Consider a preference dataset $\mathcal{D}=\{(u_{i},x_{i},y_{i}^{(w)},y_{i}^{(l)})\}_{i=1}^{N}$, where $u_{i}\in\{1..K\}$ is a user identifier, $x_{i}$ is a prompt, and $y_{i}^{(w)},y_{i}^{(l)}$ are chosen/winning and rejected/loosing completions. We partition users into $\mathcal{U}_{\mathrm{train}}$ (for learning shared preference structure) and $\mathcal{U}_{\mathrm{adapt}}$ (for evaluating few-shot personalization). Users in $\mathcal{U}_{\mathrm{adapt}}$ are further split into support sets $\mathcal{D}_{k}^{\mathrm{support}}$ (for adaptation) and query sets $\mathcal{D}_{k}^{\mathrm{query}}$ (for evaluation). For user $k$, we denote their full dataset as $\mathcal{D}_{k}=\{(x_{i},y_{i}^{(w)},y_{i}^{(l)}):u_{i}=k\}$.
Personalized Reward Modeling.  The reward model conditions on user identity: $r_{\theta,z_{k}}(y\mid x)$, decomposing into shared parameters $\theta$ (general preference structure) and user-specific parameters $z_{k}$ (individual preferences). Training on $\mathcal{U}_{\mathrm{train}}$ learns both $\theta$ and $\{z_{k}\}_{k\in\mathcal{U}_{\mathrm{train}}}$. At deployment, for user $k\in\mathcal{U}_{\mathrm{adapt}}$ with support data $\mathcal{D}_{k}^{\mathrm{support}}$, we adapt:
$z_{k}=\mathcal{A}\big(\mathcal{D}_{k}^{\mathrm{support}};\theta\big)$
where $\mathcal{A}$ is the adaptation algorithm.
@@ 3.2 Deployment via Reward-Guided Decoding
Given computational constraints prohibiting per-user policy fine-tuning, we deploy personalized alignment through inference-time guidance using Reward-Guided Decoding Khanov et al. (2024).
Reward-Guided Decoding (ARGS).  At each generation step $t$, ARGS scores the top-$k$ tokens $V^{(k)}_{t}$ retrieved from an off-the-shelf prior LLM policy, $\pi$, using
@@
$\quad+\lambda\cdot r_{\theta,z_{k}}(v\mid x,y_{<t}),$
@@ 3.3 The Evaluation Gap
Standard practice evaluates personalization methods by reward model accuracy, defined below.
@@
For user $k$ with evaluation set $\mathcal{D}_{k}^{\text{query}}=\{(x_{i},y_{i}^{(w)},y_{i}^{(l)})\}$, we define the Reward Model Ranking Accuracy as
$\frac{1}{|\mathcal{D}_{k}^{\text{eval}}|}\sum_{\mathcal{D}_{k}^{\text{eval}}}\mathbb{I}\Big[r_{\theta,z_{k}}(y_{i}^{(w)}\mid x_{i})>r_{\theta,z_{k}}(y_{i}^{(l)}\mid x_{i})\Big].$ (2)
This measures pairwise ranking on complete responses of the reward model. Later we will show that this standard metric has several issues and is not predictive of deployment performance.
4 Policy Accuracy: Measuring Preference Ranking Under RGD
We introduce a metric quantifying whether the RGD scoring function—not just the reward model in isolation—correctly ranks preferred over dispreferred responses.
@@ 4 Policy Accuracy: Measuring Preference Ranking Under RGD
Let $s:\mathcal{Y}\times\mathcal{X}\to\mathbb{R}$ be the scoring function used at generation time. The policy accuracy for user $k$ is given by
$\frac{1}{|\mathcal{D}_{k}^{\text{query}}|}\sum_{\mathcal{D}_{k}^{\text{query}}}\mathbb{I}\Big[s(y_{i}^{(w)}\mid x_{i})>s(y_{i}^{(l)}\mid x_{i})\Big],$ (3)
where $y^{(w)}$ and $y^{(l)}$ denote the chosen (winning) and rejected (losing) completions.
We instantiate $s$ with three scoring functions, each revealing different aspects of the personalization pipeline.
Base Policy. The base policy’s length-normalized log-likelihood, its off-the-shelf non-personalized zero-shot ranking ability,
$s_{\text{base}}(y\mid x)=\frac{1}{|y|}\sum_{t=1}^{|y|}\log\pi(y_{t}\mid x,y_{<t}).$ (4)
Global RGD. RGD with a non-personalized reward model $r_{\theta}$ trained by aggregating data across all users denoted $s_{\text{global}}(y\mid x)$,
$\sum_{t=1}^{|y|}\Big[\log\pi(y_{t}\mid x,y_{<t})+\lambda\cdot r_{\theta}(y_{t}\mid x,y_{<t})\Big].$ (5)
Personalized RGD. RGD with personalized reward model $r_{\theta,z_{k}}$ denoted $s_{\text{personal}}(y\mid x;z_{k})$,
$\sum_{t=1}^{|y|}\Big[\log\pi(y_{t}\mid x,y_{<t})+\lambda\cdot r_{\theta,z_{k}}(y_{t}\mid x,y_{<t})\Big].$ (6)
Comparing these reveals: (1) whether reward guidance improves ranking over the base policy, and (2) whether personalization provides gains beyond a global reward model.
In-Context Learning. An alternative personalization mechanism conditions the base policy on user-specific demonstrations $\mathcal{D}_{k}^{\text{demo}}=\{(x_{i},y_{i}^{(w)},y_{i}^{(l)}\}\subset\mathcal{D}_{k}^{\text{support}}$ rather than learning reward parameters. The ICL scoring function, denoted $s_{\text{ICL}}(y\mid x;\mathcal{D}_{k}^{\text{demo}})$ is
$\frac{1}{|y|}\sum_{t=1}^{|y|}\log\pi(y_{t}\mid\mathcal{D}_{k}^{\text{demo}},x,y_{<t}),$ (7)
[…]
Outstanding Limitation. Policy accuracy measures how well the scoring function ranks static responses—not whether the policy will actually generate outputs that actually align with user preferences. A method might rank existing responses correctly while producing generations that differ substantially from what users would write. This motivates our behavioral evaluation in Section 5.